A 1200 kg automobile travels at 9 m/s.  

  1. What is its kinetic energy? 
  2. What net work would be required to bring it to a stop? 
1
Expert's answer
2021-08-05T11:35:56-0400

Given:

"m=1200\:\rm kg"

"v=9\:\rm m/s"

1)

"E_k=\frac{mv^2}{2}=\frac{1200*9^2}{2}=48600\:\rm J"

2)

"W=-\Delta E_k=-48600\:\rm J"